OpenStudy (anonymous):

What's the answer to 4(6z-1)-5(z+9)=6(z+1)

OpenStudy (math2400):

first you nee to distribute. know how to do that?

OpenStudy (anonymous):

24z-4-5z-45=6z+6 ; 19z-6z=6+49 ; 13z= 55; z=55/13

OpenStudy (anonymous):

z=55/13